OpenStudy (anonymous):

is the equation of d over five = 4 d=20?

OpenStudy (compassionate):

Hi, \[\frac{ 20 }{ 4} = 5\] Twenty divided by four is five, not four. If, d, is set to twenty, the solution cannot be four. Also - please change your profile picture to something appropriate.

OpenStudy (anonymous):

\[\frac{ d }{ 5 }=4\]

OpenStudy (anonymous):

I Shall my friends the one who made this for me @Compassionate sorry for that

OpenStudy (anonymous):

20÷5=4

OpenStudy (compassionate):

That picture is an old caption on 4Chan. I know because I use it often.